OpenStudy (anonymous):

A culture started with 5,000 bacteria. After 4hours, it grew to 6,000 bacteria. Predict how many bacteria will be present after 17 hours. Round your answer to the nearest whole number. P=Ae^kt PLEASE HELP, IDK HOW TO DO THIS

OpenStudy (anonymous):

so, it is 1000 bact every four hours

OpenStudy (anonymous):

yes i believe so

OpenStudy (anonymous):

so 17 hours isa odd number,

OpenStudy (anonymous):

maybe 4/17 try that

OpenStudy (anonymous):

idk how to use this formula P=Ae^kt

OpenStudy (anonymous):

Not 1000 every 4 hours. Just 1000 over the first 4 hours. In the formula, P is the amount at any time, A is the initial amount, k is a constant we have to find, and t is the time.

OpenStudy (anonymous):

Start with P = 5000e^(kt)

OpenStudy (anonymous):

Plug in the given values for P and t and solve for k.

OpenStudy (anonymous):

can you please give me the answer i have 2 minutes left and i've never seen a problem like this before

OpenStudy (anonymous):

hello?

OpenStudy (triciaal):

@boogaboo this is not a linear function where the rate is constant

OpenStudy (triciaal):

P=Ae^kt

OpenStudy (triciaal):

when P = 5000 t =0 when P = 6000 t = 4 find P when t = 17

OpenStudy (anonymous):

you wanna do this the quick snappy way?

OpenStudy (anonymous):

grows by a factor of \(\frac{6000}{5000}=\frac{6}{5}\) every 4 hours model as \[\huge P(t)=5000\times( \frac{6}{5})^{\frac{t}{4}}\]

OpenStudy (anonymous):

if you want to know the population in 17 hours replace \(t\) by \(17\) and compute \[\huge P(17)=5000\times( \frac{6}{5})^{\frac{17}{4}}\]